# The X-ray properties and structure of A3571 up to R500

### v1

@@
Abell 3571 is a nearby, X-ray bright galaxy cluster located in the Shapley Supercluster. Although it appears morphologically relaxed in X-ray images, multiwavelength observations reveal subtle indications of residual dynamical activity, likely associated with past merger events. Using wide-field () data from the Einstein Probe Follow-up X-ray Telescope (EP-FXT), we extend measurements of the cluster’s properties beyond its radius. We detect surface-brightness excesses on both the northern and southern sides, consistent with previous XMM-Newton results. The temperature, pressure, and entropy in the northern excess region are lower than the average values, whereas those on the southern side are slightly higher. However, we find no evidence for cold fronts or shocks. These features can be interpreted as sloshing motions triggered by an off-center minor merger. Our findings suggest that, despite its symmetric appearance, A3571 is still recovering from a minor merger and is currently in a post-merger phase. This work also demonstrates the excellent capability of EP-FXT for probing the outskirts of galaxy clusters.
@@
X-ray observations have revealed the characteristic signatures of such dynamical activity, confirming the widespread presence of non-equilibrium structures in galaxy clusters, among which shocks and cold fronts are the most representative. Shocks, typically generated by high-velocity mergers, heat the ICM on short timescales and appear in X-ray images as sharp, arc-shaped edges(Markevitch and Vikhlinin 2007). The Bullet Cluster (Koda et al. 2008; Markevitch et al. 2002) and Abell 3667 (Owers et al. 2009) provide two of the most striking examples. Cold fronts, by contrast, appear as contact discontinuities where denser but cooler gas is juxtaposed against hotter surroundings(Markevitch et al. 2000; Vikhlinin et al. 2001; Markevitch and Vikhlinin 2007). They are generally linked to the displacement of cool core gas induced by gravitational perturbations, such as those triggered by minor mergers (Ascasibar and Markevitch 2006; Roediger et al. 2011; ZuHone et al. 2010). In relaxed clusters, such perturbations can drive sloshing motions of the central gas, producing multiple cold fronts at different radii. Over time, these sloshing motions may evolve into intricate spiral structures, which are observed as concentric edges in X-ray images (Sonkamble et al. 2024; Kadam et al. 2024a, b). To date, spiral-like cold front structures have been identified in a number of well-studied clusters, including Abell 496 (Ghizzardi et al. 2014), Abell 795 (Kadam et al. 2024a), Perseus (Walker et al. 2017; Fabian et al. 2006a), and Virgo (Roediger et al. 2011; Zheng et al. 2025).
@@
Abell 3571 (A3571) is a nearby rich galaxy cluster at redshift 0.039 (Rossetti and Molendi 2010), located in the Shapley Supercluster (Raychaudhury et al. 1991). The brightest cluster galaxy (BCG) of A3571 is MCG–05–33–002, which exhibits a pronounced north–south elongation extending over several hundred kpc (Kemp and Meaburn 1991). Although A3571 is classified as a relaxed cool-core cluster, Laganá et al. (2019) found no evidence of a central temperature drop in its temperature map. In addition, Hudaverdi et al. (2005) reported a noticeable north-south temperature asymmetry, suggesting that A3571 may have experienced recent or ongoing dynamical evolution.
@@
Early X-ray observations revealed that the ICM of A3571 exhibited an average temperature of keV from EXOSAT observations (Edge et al. 1990) and keV based on ASCA data (Markevitch et al. 1998) (the radial temperature profile extends to at ASCA’s half-power diameter (HPD)), with a central metal abundance reaching as high as 0.6 times the solar value at R < and declining to about 0.3 at a radius of , while the XMM-Newton radial profiles further trace the temperature and metallicity distributions out to approximately (Hudaverdi et al. 2005). ROSAT imaging revealed a nearly spherical X-ray brightness distribution, with the intracluster gas appearing azimuthally symmetric, exhibiting only slight ellipticity (Nevalainen et al. 2000). Based on high-resolution observations from XMM-Newton, the central region of A3571 shows pronounced temperature and metallicity inhomogeneities, which appear as a patchy distribution. A temperature difference of up to 2 keV is detected between the northern and southern regions, suggesting that the cluster may have undergone complex dynamical processes, such as the infall of subgroups (Hudaverdi et al. 2005).
@@
Studies of A3571 not only examined the properties of the hot gas but also attempted to infer the gravitational mass and dark matter distribution within the cluster. Nevalainen et al. (2000) combined ASCA temperature data with ROSAT imaging data to estimated the total mass of the cluster to be approximately with Mpc at 90% confidence, which the cluster is in hydrostatic equilibrium. Subsequently, Nevalainen et al. (2001) derived the total mass of A3571, , by combining the BeppoSAX and ASCA temperature profiles with a dark matter density model under the assumption of hydrostatic equilibrium. The corresponding gas mass fraction was estimated to be .
@@
Although X-ray observations indicate that A3571 exhibits an overall spherically symmetric structure with a central cooling flow, multiwavelength data have revealed a more complex dynamical history. Early studies found that the spatial distribution of galaxies and the pronounced north–south elongation of the central cD galaxy are prominent characteristics (Kemp and Meaburn 1991), possibly suggesting that the cluster underwent merger or disturbance events. Subsequently, spectroscopic analyses revealed an asymmetric radial velocity distribution among member galaxies and identified multiple velocity substructures, further supporting an evolutionary scenario in which several smaller galaxy groups may have successively merged into the main cluster (Quintana and De Souza 1993). Venturi et al. (2002) conducted a radio analysis of the galaxy cluster complex consisting of A3571, A3572, and A3575, and concluded that the system is in a late merger stage, with its radio properties reflecting a relatively relaxed dynamical state. More recent multiwavelength observations have not detected large-scale radio halos or significant AGN activity, indicating that A3571 is currently in a relatively relaxed, post-merger evolutionary stage.
@@
Early multiwavelength observations suggest that A3571 is in a relatively relaxed, late post-merger stage. However, our understanding of its detailed merger history remains limited, particularly regarding the hot gas in the outskirts. Therefore, we aim to conduct a more systematic and in-depth investigation of the gas distribution and merger state of A3571 using large field-of-view (FoV) X-ray imaging observations.
@@
EP-FXT is one of the key payloads aboard the Einstein Probe (EP) mission, which was launched on January 9, 2024 (Yuan et al. 2022, 2025). It features a FoV of approximately 1 square degree, an angular resolution of per pixel, and a spatial resolution of in terms of HPD, with an energy band of 0.3-10 keV. The EP-FXT also benefits from an exceptionally low particle background, making it well suited for detecting faint diffuse X-ray emission.
@@
This paper is organised as follows; in Sect. 2, we describe the data used in our analysis. In Sect. 3, the analysis of the X-ray and optical observations is described. In Sect. 4, we discuss cluster dynamics and possible subclusters. We summarise our results in Sect. 5. We adopt (Planck Collaboration et al. 2020), and the best-fit parameters from the spectral fitting are reported with 90% confidence intervals.
@@ 2.1 X-ray data
During the Performance Verification (PV) phase of the Einstein Probe mission, A3571 was observed with the EP-FXT in full-frame (FF) mode from March 30 to 31, 2024 (ObsID 11904194436), with the total effective exposure time of 40ks. The data reduction followed standard procedures using the Follow-up X-ray Telescope Data Analysis Software (FXTDAS, version 1.20)111http://epfxt.ihep.ac.cn/analysis (Zhao et al. 2025), with additional support from general analysis tools provided by CIAO (version 4.16)222https://cxc.cfa.harvard.edu/ciao/ and HEASoft (version 6.33.1)333https://heasarc.gsfc.nasa.gov/docs/software/lheasoft.
We used the fxtchain pipeline to generate a series of calibrated event files for subsequent scientific analysis. To eliminate the effect of out-of-time (OoT) events, which account for approximately 0.23% of the total detected events in the FF mode, simulated OoT event files were created using fxtootest and were subsequently subtracted from the original images using ftimgcalc. Since the particle background is uniformly distributed and not affected by the telescope’s vignetting function, it was removed before applying vignetting correction. The background spectrum was extracted from the storage area of the event file and modeled using fxtbkggen (Zhang et al. 2025), after which it was subtracted from the image. Vignetting, caused by the decrease in effective area at larger off-axis angles, leads to reduced photon flux at the edges of the FoV. To correct for this effect and improve the SNR, we used reproject_image from CIAO to co-align and merge the images and vignetting-corrected exposure maps from FXT-A and FXT-B. The merged images were then divided by the combined exposure map using farith to produce a vignetting-corrected image. After vignetting correction, point sources were identified using the wavdetect from CIAO and excluded to avoid contamination in the analysis of diffuse emission. Subsequently, the smooth were conducted with a Gaussian kernel. The final X-ray image in the 0.3–7.0 keV band is presented in Fig. 1.
To compare EP-FXT with XMM-Newton, we selected the EPIC-MOS1 observation of A3571 conducted by XMM-Newton in July 2002 (ObsID 0086950201, with an exposure time of 33 ks) and processed the data using the XMM-Newton Science Analysis System (SAS 444https://www.cosmos.esa.int/web/xmm-newton/sas) in FF mode. Time intervals affected by flares were removed by screening the high-energy light curve.
@@
The photometric catalog used in this study was derived from the DESI Legacy Imaging Surveys DR10 (Dey et al. 2019). We select a field encompassing the A3571 cluster and its surrounding region, spanning and , with a magnitude limit of magr 22 . Cluster member candidates are identified using the red sequence method, by selecting galaxies within 3 of the best-fit red sequence and further restricting their photometric redshifts to the range . This yields a sample of 399 galaxies.
To examine the spatial distribution of galaxies in and around A3571, we generate a smoothed luminosity density map based on this photometric sample. Following the method of Wen and Han (2013), the r-band luminosity of each galaxy is estimated using its redshift and apparent magnitude. These luminosities are then mapped onto a 200 200 grid covering the selected field, and a two-dimensional (2D) Gaussian kernel with a full width of is applied to smooth the distribution.
@@
There is another cluster A3572 located about to the south of A3571. Struble and Rood (1999) estimated a redshift of 0.052 and a velocity dispersion of 1216 km s-1 with 6 galaxies. While Lauer et al. (2014) measured a redshift of based on an elliptical galaxy thought to be its BCG. Based on the photometric luminosity density map and the corresponding X-ray emission distribution, no significant enhancement in galaxy number density or X-ray surface brightness is observed at the location of A3572. Thus we conclude that A3572 is not a real cluster.
@@ 3 ANALYSIS and RESULTS
The galaxy density distribution of A3571 shows a noticeable elongation along the north–south direction, suggesting that the cluster may not be in a state of hydrostatic equilibrium. To investigate its thermodynamic properties and assess potential merger activity, we conducted a detailed imaging and spectral analysis of the ICM.
@@ 3.1.1 The Surface brightness profile
Black points are the EP-FXT data, which extends out to approximately . The blue horizontal line represents the EP-FXT particle background cts/s/arcmin2 , generated using fxtbkggen as described by Zhang et al. (2022). The green horizontal line represents the EP-FXT sky background , obtained by fitting a constant to the surface brightness in the – region of A3571, which nevertheless includes faint emission from the cluster. The blue points indicate the EP-FXT surface brightness after subtracting both background components. We employed a double model to fit the background-subtracted surface brightness profile. The model equation is given below:
The best-fit parameters are listed in Table 1. Additionally, we plotted the two components of the double model, represented by the green and orange dashed lines, which reveal the presence of a dense core in the central region of A3571.
Similarly, we performed a radial analysis of the XMM-Newton MOS1 data and normalized its surface brightness profile to the maximum value of the EP-FXT data. The background components included the sky background obtained from blank-sky observations, and the particle background derived from the 10 – 12 keV spectrum. The cyan points represent the normalized XMM-Newton data, while the blue and green dashed lines correspond to the normalized particle background and the normalized sky background , respectively.
@@
Note. Best-fit parameters of the double model with errors, based on the background-subtracted. and are in units of arcmin. A, B and C are in units of .
@@ 3.1.2 Global spectral analysis
The particle background spectra of both FXT-A and FXT-B exhibit similar shapes, remaining at a level of approximately counts s-1 keV-1, which is well below the source emission. Three prominent instrumental lines are present: Al-K (1.48 keV), Fe-K (6.38 keV), and Cu-K (8.04 keV). Around 1 keV, where the source emission peaks, the particle background contributes less than 1% of the source count rate, thereby enabling more reliable analysis of faint diffuse sources. Below 1 keV, the FXT-B spectrum differs from that of FXT-A, primarily due to FXT-B’s larger effective area in this energy band. In addition, extended sources are influenced by atmospheric absorption at low energies because of the satellite’s low-Earth orbit. Therefore, we restrict our data processing and analysis to the 1.0–8.0 keV range.

### v2

@@
Abell 3571 is a nearby, X-ray-bright galaxy cluster located in the Shapley Supercluster. Although it appears morphologically relaxed in X-ray images, multiwavelength observations reveal subtle indications of residual dynamical activity likely associated with past merger events. Using wide-field () data from the Einstein Probe Follow-up X-ray Telescope (EP-FXT), we extended measurements of the cluster’s properties beyond its radius. We detect surface-brightness excesses on both the northern and southern sides, consistent with previous XMM-Newton results. The temperature, pressure, and entropy in the northern excess region are lower than the average values, whereas those on the southern side are slightly higher. However, we find no evidence of cold fronts or shocks. These features can be interpreted as sloshing motions triggered by an off-center minor merger. Our findings suggest that, despite its symmetric appearance, A3571 is still recovering from a minor merger and is currently in a post-merger phase. This work also demonstrates the excellent capability of EP-FXT for probing the outskirts of galaxy clusters.
@@
X-ray observations have revealed the characteristic signatures of such dynamical activity, confirming the widespread presence of nonequilibrium structures in galaxy clusters, among which shocks and cold fronts are the most representative. Shocks, typically generated by high-velocity mergers, heat the ICM on short timescales and appear in X-ray images as sharp, arc-shaped edges (Markevitch and Vikhlinin 2007). The Bullet Cluster (Koda et al. 2008; Markevitch et al. 2002) and Abell 3667 (Owers et al. 2009) provide two of the most striking examples. Cold fronts, by contrast, appear as contact discontinuities where denser but cooler gas is juxtaposed against hotter surroundings (Markevitch et al. 2000; Vikhlinin et al. 2001; Markevitch and Vikhlinin 2007). They are generally linked to the displacement of cool core gas induced by gravitational perturbations, such as those triggered by minor mergers (Ascasibar and Markevitch 2006; Roediger et al. 2011; ZuHone et al. 2010). In relaxed clusters, such perturbations can drive sloshing motions of the central gas, producing multiple cold fronts at different radii. Over time, these sloshing motions can evolve into intricate spiral structures, which are observed as concentric edges in X-ray images (Sonkamble et al. 2024; Kadam et al. 2024a, b). To date, spiral-like cold front structures have been identified in a number of well-studied clusters, including Abell 496 (Ghizzardi et al. 2014), Abell 795 (Kadam et al. 2024a), Perseus (Walker et al. 2017; Fabian et al. 2006a), and Virgo (Roediger et al. 2011; Zheng et al. 2025).
@@
Abell 3571 (A3571) is a nearby rich galaxy cluster at redshift 0.039 (Rossetti and Molendi 2010), located in the Shapley Supercluster (Raychaudhury et al. 1991). The brightest cluster galaxy (BCG) of A3571 is MCG–05–33–002, which exhibits a pronounced north–south elongation extending over several hundred kiloparsecs (Kemp and Meaburn 1991). Although A3571 is classified as a relaxed cool-core cluster, Laganá et al. (2019) found no evidence of a central temperature drop in its temperature map. In addition, Hudaverdi et al. (2005) reported a noticeable north–south temperature asymmetry, suggesting that A3571 recently experienced – or is still experiencing – dynamical evolution.
@@
Early X-ray observations revealed that the ICM of A3571 exhibited an average temperature of keV, based on EXOSAT111The European X-ray Observatory Satellitedata (Edge et al. 1990), or keV, based on ASCA222The Advanced Satellite for Cosmology and Astrophysics data (the radial temperature profile extends to at ASCA’s half-power diameter; Markevitch et al. 1998), with a central metal abundance reaching as high as 0.6 times the solar value at R < and declining to about 0.3 at a radius of . XMM-Newton radial profiles further trace the temperature and metallicity distributions out to approximately (Hudaverdi et al. 2005). ROSAT333The ROentgen SATellite imaging revealed a nearly spherical X-ray brightness distribution, with the intracluster gas appearing azimuthally symmetric and exhibiting only slight ellipticity (Nevalainen et al. 2000). Based on high-resolution observations from XMM-Newton, the central region of A3571 shows pronounced temperature and metallicity inhomogeneities, which appear as a patchy distribution. A temperature difference of up to 2 keV is detected between the northern and southern regions, suggesting that the cluster has undergone complex dynamical processes, such as the infall of subgroups (Hudaverdi et al. 2005).
@@
Studies of A3571 not only examined the properties of the hot gas but also attempted to infer the gravitational mass and dark matter distribution within the cluster. Nevalainen et al. (2000) combined ASCA temperature data with ROSAT imaging data to estimated the total mass of the cluster to be approximately with Mpc at 90% confidence, which means the cluster is in hydrostatic equilibrium. Subsequently, Nevalainen et al. (2001) derived the total mass of A3571, , by combining the BeppoSAX and ASCA temperature profiles with a dark matter density model under the assumption of hydrostatic equilibrium. The corresponding gas mass fraction was estimated to be .
@@
Although X-ray observations indicate that A3571 exhibits an overall spherically symmetric structure with a central cooling flow, multiwavelength data have revealed a more complex dynamical history. Early studies found that the spatial distribution of galaxies and the pronounced north–south elongation of the central cD galaxy are prominent characteristics (Kemp and Meaburn 1991), possibly suggesting that the cluster underwent merger or disturbance events. Subsequently, spectroscopic analyses revealed an asymmetric radial velocity distribution among member galaxies and identified multiple velocity substructures, further supporting an evolutionary scenario in which several smaller galaxy groups successively merged into the main cluster (Quintana and De Souza 1993). Venturi et al. (2002) conducted a radio analysis of the galaxy cluster complex consisting of A3571, A3572, and A3575, and concluded that the system is in a late merger stage, with its radio properties reflecting a relatively relaxed dynamical state. More recent multiwavelength observations have not detected large-scale radio halos or significant active galactic nucleus activity, indicating that A3571 is currently in a relatively relaxed, post-merger evolutionary stage.
@@
Early multiwavelength observations suggest that A3571 is in a relatively relaxed, late post-merger stage. However, our understanding of its merger history remains limited, particularly regarding the hot gas in the outskirts. Therefore, we aim to conduct a more systematic and in-depth investigation of the gas distribution and merger state of A3571 using large field-of-view (FoV) X-ray imaging observations.
@@
The Follow-up X-ray Telescope (FXT) is one of the key payloads aboard the Einstein Probe (EP) mission, which was launched on January 9, 2024 (Yuan et al. 2022, 2025). It features a FoV of approximately 1 square degree, an angular resolution of per pixel, and a half-power diameter spatial resolution of , with an energy band of 0.3-10 keV. The EP-FXT also benefits from an exceptionally low particle background, making it well suited for detecting faint diffuse X-ray emission.
@@
This paper is organized as follows: In Sect. 2 we describe the data used in our analysis. In Sect. 3 the analysis of the X-ray and optical observations is described. In Sect. 4 we discuss cluster dynamics and possible subclusters. We summarize our results in Sect. 5. We adopt and (Planck Collaboration et al. 2020), and the best-fit parameters from the spectral fitting are reported with 90% confidence intervals.
@@ 2.1 X-ray data
During the performance verification phase of the EP mission, A3571 was observed with the EP-FXT in full-frame (FF) mode from March 30 to 31, 2024 (ObsID 11904194436), with the total effective exposure time of 40ks. The data reduction followed standard procedures using the FXT Data Analysis Software (FXTDAS, version 1.20)444http://epfxt.ihep.ac.cn/analysis (Zhao et al. 2025), with additional support from general analysis tools provided by CIAO (version 4.16)555https://cxc.cfa.harvard.edu/ciao/ and HEASoft (version 6.33.1)666https://heasarc.gsfc.nasa.gov/docs/software/lheasoft.
We used the fxtchain pipeline to generate a series of calibrated event files for subsequent scientific analysis. To eliminate the effect of out-of-time events, which account for approximately 0.23% of the total detected events in the FF mode, simulated out-of-time event files were created using fxtootest and were subsequently subtracted from the original images using ftimgcalc. Since the particle background is uniformly distributed and not affected by the telescope’s vignetting function, it was removed before applying vignetting correction. The background spectrum was extracted from the storage area of the event file and modeled using fxtbkggen (Zhang et al. 2025), after which it was subtracted from the image. Vignetting, caused by the decrease in effective area at larger off-axis angles, leads to reduced photon flux at the edges of the FoV. To correct for this effect and improve the S/N, we used reproject_image from CIAO to co-align and merge the images and vignetting-corrected exposure maps from FXT-A and FXT-B. The merged images were then divided by the combined exposure map using farith to produce a vignetting-corrected image. After vignetting correction, point sources were identified using the wavdetect from CIAO and excluded to avoid contamination in the analysis of diffuse emission. Subsequently, the smooth were conducted with a Gaussian kernel. The final X-ray image in the 0.3–7.0 keV band is presented in Fig. 1.
To compare EP-FXT with XMM-Newton, we selected the European Photon Imaging Camera (EPIC) MOS1 observation of A3571 conducted by XMM-Newton in July 2002 (ObsID 0086950201, with an exposure time of 33 ks) and processed the data using the XMM-Newton Science Analysis System (SAS777https://www.cosmos.esa.int/web/xmm-newton/sas) in FF mode. Time intervals affected by flares were removed by screening the high-energy light curve.
@@ 2.2 Optical data
The photometric catalog used in this study was derived from the Dark Energy Spectroscopic Instrument (DESI) Legacy Imaging Surveys Data Release 10 (Dey et al. 2019). We selected a field encompassing the A3571 cluster and its surrounding region, spanning and , with a magnitude limit of magr 22 . Cluster member candidates are identified using the red sequence method, by selecting galaxies within 3 of the best-fit red sequence and further restricting their photometric redshifts to the range . This yields a sample of 399 galaxies.
To examine the spatial distribution of galaxies in and around A3571, we generated a smoothed luminosity density map based on this photometric sample. Following the method of Wen and Han (2013), the r-band luminosity of each galaxy was estimated using its redshift and apparent magnitude. These luminosities were then mapped onto a 200 200 grid covering the selected field, and a 2D Gaussian kernel with a full width of was applied to smooth the distribution.
@@
There is another galaxy cluster candidate, A3572, located about to the south of A3571. Struble and Rood (1999) estimated a redshift of 0.052 and a velocity dispersion of 1216 km s-1 based on six galaxies, while Lauer et al. (2014) reported a redshift of for an elliptical galaxy that is considered to be its BCG. Based on the photometric luminosity density map and the corresponding X-ray emission distribution, we do not detect any significant enhancement in galaxy number density or X-ray surface brightness at the reported location of A3572. This suggests that, if A3572 exists, it may host a very weak ICM or be a poor system that is not clearly detectable with the currently available photometric and X-ray data. Consequently, the nature of A3572 remains uncertain.
@@ 3 Analysis and results
The galaxy density distribution of A3571 shows a noticeable elongation along the north–south direction, suggesting that the cluster is not in a state of hydrostatic equilibrium. To investigate its thermodynamic properties and assess potential merger activity, we conducted a detailed imaging and spectral analysis of the ICM.
@@ 3.1.1 The surface brightness profile
Black points are the EP-FXT data, which extends out to approximately . The horizontal blue line represents the EP-FXT particle background cts/s/arcmin2 , generated using fxtbkggen as described by Zhang et al. (2022). The horizontal green line represents the EP-FXT sky background , obtained by fitting a constant to the surface brightness in the – region of A3571, which nevertheless includes faint emission from the cluster. The blue points indicate the EP-FXT surface brightness after subtracting both background components. We employed a double- model to fit the background-subtracted surface brightness profile. The model equation is
The best-fit parameters are listed in Table 1. Additionally, we plot in Fig. 3 the two components of the double- model, which reveal the presence of a dense core in the central region of A3571.
Similarly, we performed a radial analysis of the XMM-Newton MOS1 data and normalized its surface brightness profile to the maximum value of the EP-FXT data. The background components included the sky background obtained from blank-sky observations, and the particle background derived from the 10 – 12 keV spectrum. The dashed blue and green lines correspond to the normalized particle background of and the normalized sky background of , respectively.
@@ 3.1.2 Global spectral analysis
The particle background spectra of FXT-A and FXT-B exhibit similar shapes, remaining at a level of approximately counts s-1 keV-1, which is well below the source emission. Three prominent instrumental lines are present: Al-K (1.48 keV), Fe-K (6.38 keV), and Cu-K (8.04 keV). Around 1 keV, where the source emission peaks, the particle background contributes less than 1% of the source count rate, thereby enabling more reliable analysis of faint diffuse sources. Below 1 keV, the FXT-B spectrum differs from that of FXT-A, primarily due to FXT-B’s larger effective area in this energy band. In addition, extended sources are influenced by atmospheric absorption at low energies because of the satellite’s low-Earth orbit. Therefore, we restricted our data processing and analysis to the 1.0–8.0 keV range.